### v2

# On Sum-Free Functions

Abstract.
A function from $\mathbb{F}_{2^{n}}$ to $\mathbb{F}_{2^{n}}$ is said to be $k$th order sum-free if the sum of its values over each $k$-dimensional $\mathbb{F}_{2}$-affine subspace of $\mathbb{F}_{2^{n}}$ is nonzero. This notion was recently introduced by C. Carlet as, among other things, a generalization of APN functions. At the center of this new topic is a conjecture about the sum-freedom of the multiplicative inverse function $f_{\text{\rm inv}}(x)=x^{-1}$ (with $0^{-1}$ defined to be 0). It is known that $f_{\text{\rm inv}}$ is 2nd order (equivalently, (n2)(n-2)th order) sum-free if and only if $n$ is odd, and it is conjectured that for $3\leq k\leq n-3$, $f_{\text{\rm inv}}$ is never $k$th order sum-free. The conjecture has been confirmed for even $n$ but remains open for odd $n$. In the present paper, we show that the conjecture holds under each of the following conditions: (1) $n=13$; (2) $3\mid n$; (3) $5\mid n$; (4) the smallest prime divisor $l$ of $n$ satisfies $(l-1)(l+2)\leq(n+1)/2$. We also determine the “right” $q$-ary generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$ in the context of sum-freedom. This $q$-ary generalization not only maintains most results for its binary version, but also exhibits some extraordinary phenomena that are not observed in the binary case.
Key words and phrases:
APN function, finite field, multiplicative inverse function, sum-free function
2010 Mathematics Subject Classification:
11G25, 11T06, 11T71, 94D10
1. Introduction
In a recent paper [4], Carlet introduced the notion of sum-free functions. A function $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional affine subspaces $A$ of $\mathbb{F}_{2^{n}}$. Unless specified otherwise, subspaces and affine subspaces of $\mathbb{F}_{2^{n}}$ are over $\mathbb{F}_{2}$. This definition does not involve the multiplicative structure of $\mathbb{F}_{2^{n}}$, hence sum-freedom can be defined for functions from $\mathbb{F}_{2}^{n}$ to $\mathbb{F}_{2}^{n}$. However, the multiplicative structure of $\mathbb{F}_{2^{n}}$ plays an important role in the constructions and analysis of such functions. As described in [4], the motivation for the notion of sum-free functions was two-fold: First, sum-free functions are a natural generalization of almost perfect nonlinear (APN) functions; the latter have been extensively studied in cryptography for their resistance against differential attacks [1, 3, 6, 19, 20, 21]. In particular, APN functions are precisely 2nd order sum-free functions [11]. Second, sum-freedom addresses the vulnerability of block ciphers to integral attacks which exploit the predicability of the sums of values of S-boxes on affine subspaces [14, 22]. Moreover, from a theoretic point of view, sum-freedom is a notion with rich mathematical content; beneath its simple appearance, there are difficult and enticing questions that require sophisticated mathematical tools.
Let $f_{\text{\rm inv}}$ denote the multiplicative inverse function of $\mathbb{F}_{2^{n}}$, that is, $f_{\text{\rm inv}}(x)=x^{-1}$ for $x\in\mathbb{F}_{2^{n}}^{*}$ and $f_{\text{\rm inv}}(0)=0$. For $n\geq 2$, we have $f_{\text{\rm inv}}(x)=x^{2^{n}-2}$ for all $x\in\mathbb{F}_{2^{n}}$. The sum-freedom of $f_{\text{\rm inv}}$ has been the focus of several recent papers [4, 5, 7]. It is well known that $f_{\text{\rm inv}}$ is APN (i.e., 2nd order sum-free) if and only if $n$ is odd [20]. Carlet [4] determined that $\sum_{x\in A}1/x\neq 0$ for all affine subspaces $A$ of $\mathbb{F}_{2^{n}}$ not containing 0. Hence $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if $\sum_{0\neq u\in E}1/u\neq 0$ for all $k$-dimensional subspaces $E$ of $\mathbb{F}_{2^{n}}$. It is trivial that $f_{\text{\rm inv}}$ is 11st order sum-free and it is known from [5] that for $1\leq k\leq n-1$, $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if it is (nk)(n-k)th order sum-free. Therefore, for $1\leq k\leq n-1$, $f_{\text{\rm inv}}$ is $k$th order sum-free if $k\in\{1,n-1\}$ for even $n$ and $k\in\{1,2,n-2,n-1\}$ for odd $n$. Theoretic results and computer searches, which we will summarize in the next section, strongly suggest that these are the only values of $k$ for which $f_{\text{\rm inv}}$ is $k$th order sum-free. This claim has been confirmed in [7] for even $n$, but for odd $n$ the question remains open. (We remind the reader that the even $n$ case does not represent “half” of the question as we will see later that a small prime divisor of $n$ gives an advantage for the treatment of the question.) For $n\geq 2$, define
(1.1) $\mathcal{K}_{n}=\{1\leq k\leq n-1:\text{$f_{\text{\rm inv}}$ is not $k$th order sum-free on $\mathbb{F}_{2^{n}}$}\}.$
(Note that the meaning of $\mathcal{K}_{n}$ here is different from that in [7].) Trivially, $\mathcal{K}_{2},\mathcal{K}_{3},\mathcal{K}_{5}$ are empty, and we already know that $\mathcal{K}_{n}=\{2,3,\dots n-2\}$ for even $n\geq 4$ and $\mathcal{K}_{n}\subset\{3,4,\dots,n-3\}$ for odd $n\geq 7$. Thus the above open question can be formulated as the following
Conjecture 1.1.
For odd $n\geq 7$, $\mathcal{K}_{n}=\{3,4,\dots,n-3\}$.
In the present paper, we obtain several new results concerning the above conjecture. We show that the conjecture holds under any of the following conditions:
  • $n=13$;
  • $3\mid n$;
  • $5\mid n$;
  • the smallest prime divisor $l$ of $n$ satisfies $(l-1)(l+2)\leq(n+1)/2$.
Because of the last condition, the most difficult cases of Conjecture 1.1 are the scenarios where $n$ is a prime or a product of two primes, not necessarily different, close to each other. Carlet [5] proved that $4\in\mathcal{K}_{n}$ for all $n\geq 6$. We found a different proof of this fact using the Lang-Weil bound. The new proof is included in the present paper because we believe that the approach can be applied to more general situations.
The notion of sum-freedom can be generalized to functions from $\mathbb{F}_{q^{n}}$ to $\mathbb{F}_{q^{n}}$, mutatis mutandis, where $q$ is a prime power. A function $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional $\mathbb{F}_{q}$-affine subspaces $A$ of $\mathbb{F}_{q^{n}}$. In this context, an immediate question is this: What is the $q$-ary version of the binary multiplicative inverse function $f_{\text{\rm inv}}$ that maintains the properties of $f_{\text{\rm inv}}$? We determine that the “right” $q$-ary version is the function $g_{q-1}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ defined by $g_{q-1}(x)=1/x^{q-1}$ with $1/0$ defined as 0. The reason for this, as we shall see in Section 4, is not entirely obvious. It turns out that most existing results concerning the sum-freedom of $f_{\text{\rm inv}}$ still hold for $g_{q-1}$. However, more interesting things seem to be happening in the $q$-ary case than the binary case. For example, through a computer search, we find that with $q=3,5$ and $n=7$, $g_{q-1}$ on $\mathbb{F}_{q^{7}}$ is $k$th order sum-free for all $1\leq k\leq 6$; similar phenomena have not been observed in the binary case. We are not aware of any immediate applications of the functions $g_{2}(x)=1/x^{2}$ on $\mathbb{F}_{3^{7}}$ and $g_{4}(x)=1/x^{4}$ on $\mathbb{F}_{5^{7}}$. Nevertheless, these are extraordinary examples from a mathematical point of view.
The paper is organized as follows: Section 2 provides the preliminaries for the paper. We recall the existing results about binary sum-free functions, especially the multiplicative inverse function $f_{\text{\rm inv}}$. We also include a determination of all (n1)(n-1)th order sum-free functions on $\mathbb{F}_{2^{n}}$; this question has not been addressed previously. In Section 3, we confirm Conjecture 1.1 under each of the conditions described above and we prove that $4\in\mathcal{K}_{n}$ for all $n\geq 6$. In Section 4, we explore $q$-ary sum-free functions with focus on the generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$. It follows from some nontrivial computation that the “right” $q$-ary version of $f_{\text{\rm inv}}$ is the function $g_{q-1}$ defined by $g_{q-1}(x)=1/x^{q-1}$. We show that most existing results concerning the sum-freedom of $f_{\text{\rm inv}}$ still hold for $g_{q-1}$. We also include some limited computer search results on $g_{q-1}$ with $q=3,5$. The appendix contains a proof of the absolute irreducibility of a symmetric homogeneous polynomial of degree 8 in 4 variables over $\mathbb{F}_{2}$. The proof is rather technical and the result is used in Section 3.
Throughout the paper, a number of tools are used in the proofs and discussions (e.g., the Moore determinant, Carlitz’s formula, symmetric polynomials, the Lang-Weil bound). The relations with other areas of mathematics is part of the reason that we find the notion of sum-free functions and the questions that arise from it attractive.
2. Preliminaries
2.1. $k$th order sum-free functions with k=0,1,2,n1,nk=0,1,2,n-1,n
Let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a function. We know that
  • $f$ is 0th order sum-free if and only if $f(x)\neq 0$ for all $x\in\mathbb{F}_{2^{n}}$;
  • $f$ is 11st order sum-free if and only if $f$ is a bijection;
  • $f$ is 22nd order sum-free if and only if $f$ is APN;
  • $f$ is $n$th order sum-free if and only if $\sum_{x\in\mathbb{F}_{2^{n}}}f(x)\neq 0$, i.e., $\deg f=2^{n}-1$ when $f$ is represented by a polynomial modulo $X^{2^{n}}-X$.
The last statement follows from the fact that
$\sum_{x\in\mathbb{F}_{2^{n}}}x^{i}=\begin{cases}0&\text{if}\ 0\leq i\leq 2^{n}-2,\cr 1&\text{if}\ i=2^{n}-1.\end{cases}$
There does not appear to be a determination of (n1)(n-1)th order sum-free functions in the literature. Hence the following theorem is useful for the record.
Theorem 2.1.
Let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be represented by
$f(X)=\sum_{i=0}^{2^{n}-1}a_{i}X^{i}\in\mathbb{F}_{2^{n}}[X].$
Then $f$ is (n1)(n-1)th order sum-free if and only if $a_{2^{n}-1}=0$ and
(2.1) $\left|\begin{matrix}a_{2^{n}-1-2^{0}}&a_{2^{n}-1-2^{1}}&\cdots&a_{2^{n}-1-2^{n-1}}\cr a_{2^{n}-1-2^{1}}^{2}&a_{2^{n}-1-2^{2}}^{2}&\cdots&a_{2^{n}-1-2^{0}}^{2}\cr\vdots&\vdots&&\vdots\cr a_{2^{n}-1-2^{n-1}}^{2^{n-1}}&a_{2^{n}-1-2^{0}}^{2^{n-1}}&\cdots&a_{2^{n}-1-2^{n-2}}^{2^{n-1}}\end{matrix}\right|\neq 0.$
Proof.
When $n=1$, the claim is obviously true, so we assume $n\geq 2$.
$(\Leftarrow$) For $0\neq b\in\mathbb{F}_{2^{n}}$, let
$A_{b}=\{x\in\mathbb{F}_{2^{n}}:\text{\rm Tr}(bx)=0\},$
$B_{b}=\{x\in\mathbb{F}_{2^{n}}:\text{\rm Tr}(bx)=1\},$
where $\text{\rm Tr}=\text{\rm Tr}_{2^{n}/2}$. Each hyperplane (affine subspace of co-dimension one) of $\mathbb{F}_{2^{n}}$ is of the form $A_{b}$ or $B_{b}$, depending on whether it contains 0 or not. For $0\neq b\in\mathbb{F}_{2^{n}}$, we have
$\sum_{x\in B_{b}}f(x)\,$ $=\sum_{\begin{subarray}{c}x\in\mathbb{F}_{2^{n}}\cr\text{\rm Tr}(bx)=1\end{subarray}}f(x)=\sum_{x\in\mathbb{F}_{2^{n}}}f(x)\text{\rm Tr}(bx)$
$=\sum_{x\in\mathbb{F}_{2^{n}}}\Bigl(\sum_{i=0}^{2^{n}-1}a_{i}x^{i}\Big)\Bigl(\sum_{j=0}^{n-1}b^{2^{j}}x^{2^{j}}\Bigr)$
$=\sum_{i=0}^{2^{n}-1}\sum_{j=0}^{n-1}a_{i}b^{2^{j}}\sum_{x\in\mathbb{F}_{2^{n}}}x^{i+2^{j}}.$
In the above,
$\sum_{x\in\mathbb{F}_{2^{n}}}x^{i+2^{j}}=\begin{cases}-1&\text{if}\ i+2^{j}=2^{n}-1,\cr 0&\text{otherwise}.\end{cases}$
Hence
$\sum_{x\in B_{b}}f(x)=\sum_{j=0}^{n-1}a_{2^{n}-1-2^{j}}b^{2^{j}}.$
By [13, Theorem 2.29], the $\mathbb{F}_{2}$-linear map $\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$, $x\mapsto\sum_{j=0}^{n-1}a_{2^{n}-1-2^{j}}x^{2^{j}}$ is bijective if and if condition (2.1) is satisfied. Therefore, $\sum_{x\in B_{b}}f(x)\neq 0$. Since $a_{2^{n}-1}=0$, we have $\sum_{x\in\mathbb{F}_{2^{n}}}f(x)=0$, hence $\sum_{x\in A_{b}}f(x)=\sum_{x\in B_{b}}f(x)\neq 0$.
($\Rightarrow$) Simply trace back the proof of ($\Leftarrow$). ∎
2.2. Known results about $f_{\text{\rm inv}}$
References [4, 5, 7] contain many results concerning the sum-freedom of the multiplicative inverse function $f_{\text{\rm inv}}$ of $\mathbb{F}_{2^{n}}$. We gather them in their strongest versions for the reader’s convenience. We assume that $n\geq 2$ and $1\leq k\leq n-1$.
  1. (1)
    If $\gcd(k,n)>1$, then $k\in\mathcal{K}_{n}$ (obvious).
  2. (2)
    $f_{\text{\rm inv}}$ is 1st order sum-free (obvious).
  3. (3)
    $f_{\text{\rm inv}}$ is 2nd order sum-free if and only if $n$ is odd ([20]).
  4. (4)
    $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if it is (nk)(n-k)th order sum-free ([5, Theorem 3]).
  5. (5)
    If $k,l\in\mathcal{K}_{n}$ and $kl<n$, then $k+l\in\mathcal{K}_{n}$ ([5, Theorem 4]).
  6. (6)
    For even $n$, $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if $k\in\{1,n-1\}$ ([7, Theorem 5.2]).
  7. (7)
    Conjecture 1.1 holds for $n=7,9,11$ ([5, Table 1]).
  8. (8)
    For $n\geq 6$, $3\in\mathcal{K}_{n}$ ([5, Theorem 6]).
  9. (9)
    For $n\geq 6$, $4\in\mathcal{K}_{n}$ ([5, Theorem 7]).
  10. (10)
    If $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{2}[X]$, then $k\in\mathcal{K}_{n}$ ([7, Corollary 3.2]).
  11. (11)
    If $n\geq 13k/3+3$, then $k\in\mathcal{K}_{n}$ ([7, Theorem 4.2]).
2.3. The Moore determinant
The Moore determinant (see [18]) plays an important role in the study of the sum-freedom of $f_{\text{\rm inv}}$ and its $q$-ary version. Let $X_{1},\dots,X_{k}$ be independent indeterminates. The Moore determinant in $X_{1},\dots,X_{k}$ over $\mathbb{F}_{q}$ is defined to be
$\Delta(X_{1},\dots,X_{k})=\left|\begin{matrix}X_{1}&\cdots&X_{k}\cr X_{1}^{q}&\cdots&X_{k}^{q}\cr\vdots&&\vdots\cr X_{1}^{q^{k-1}}&\cdots&X_{k}^{q^{k-1}}\end{matrix}\right|\in\mathbb{F}_{q}[X_{1},\dots,X_{k}].$
Lemma 2.2 ([16, Lemma 3.51]).
We have
$\Delta(X_{1},\dots,X_{k})=\prod_{i=1}^{k}\prod_{a_{1},\dots,a_{i-1}\in\mathbb{F}_{q}}\Bigl(X_{i}-\sum_{j=1}^{i-1}a_{j}X_{j}\Bigr).$
Corollary 2.3.
A set of $k$ elements $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ are linearly independent over $\mathbb{F}_{q}$ if and only if $\Delta(v_{1},\dots,v_{k})\neq 0$.
For $0\leq i\leq k$, let
$\Delta_{i}(X_{1},\dots,X_{k})=\left|\begin{matrix}X_{1}&\cdots&X_{k}\cr\vdots&&\vdots\cr X_{1}^{q^{i-1}}&\cdots&X_{k}^{q^{i-1}}\cr X_{1}^{q^{i+1}}&\cdots&X_{k}^{q^{i+1}}\cr\vdots&&\vdots\cr X_{1}^{q^{k}}&\cdots&X_{k}^{q^{k}}\end{matrix}\right|.$
Note that $\Delta_{k}(X_{1},\dots,X_{k})=\Delta(X_{1},\dots,X_{k})$ and $\Delta_{0}(X_{1},\dots,X_{k})=\Delta(X_{1},\dots,X_{k})^{q}$.
2.4. Criteria for $f_{\text{\rm inv}}$ not to be $k$th order sum-free
Recall that $k\in\mathcal{K}_{n}$ means that $f_{\text{\rm inv}}$ is not $k$th order sum-free. The following criterion was the basis of an approach in [7]. We follow the notation of §2.3 with $q=2$.
Theorem 2.4 ([5, Proposition 1], [7, §3.1]).
We have $k\in\mathcal{K}_{n}$ if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{2^{n}}$ such that $\Delta_{1}(v_{1},\dots,v_{k})=0$ but $\Delta(v_{1},\dots,v_{k})\neq 0$.
In the above, $\Delta(v_{1},\dots,v_{k})\neq 0$ if and only if $v_{1},\dots,v_{k}$ are linearly independent over $\mathbb{F}_{2}$.
There is another criterion by Carlet [5] for $k\in\mathcal{K}_{n}$. We will revisit this criterion in Section 3.
3. New Results on Conjecture 1.1
Recall that the set $\mathcal{K}_{n}$ is defined in equation (1.1) and Conjecture 1.1 states that for odd $n\geq 7$, $\mathcal{K}_{n}=\{3,4,\dots,n-3\}$. $n=13$ is the first value for which Conjecture 1.1 was not confirmed previously. We now confirm this case with a computer search.
Theorem 3.1.
Conjecture 1.1 is true for $n=13$.
Proof.
By §2.2 Results (4) and (8), we only have to show that $4,5,6\in\mathcal{K}_{13}$. To this end, for $k=4,5,6$, we search for $v_{1},\dots,v_{k}\in\mathbb{F}_{2^{13}}$, linearly independent over $\mathbb{F}_{2}$, such that $\Delta_{1}(v_{1},\dots,v_{k})=0$. Obviously we may assume $v_{1}=1$ (since $f_{\text{\rm inv}}$ sums to 0 on $\langle v_{1},\dots,v_{k}\rangle$, the subspace spanned by $v_{1},\dots,v_{k}$, if and only if it sums to 0 on $\langle av_{1},\dots,av_{k}\rangle$ for any $a\in\mathbb{F}_{2^{n}}^{*}$.) We represent $\mathbb{F}_{2^{13}}$ as $\mathbb{F}_{2}[X]/(f)$, where $f=X^{13}+X^{12}+X^{11}+X^{8}+1\in\mathbb{F}_{2}[X]$ is irreducible (in fact, primitive). The following are examples of the search results:
$k=4$.
$\begin{cases}v_{2}=X+X^{2}+X^{5}+X^{8}+X^{11},\cr v_{3}=X^{4}+X^{5}+X^{7}+X^{8}+X^{11},\cr v_{4}=1+X^{2}+X^{5}+X^{8}+X^{10}+X^{11}.\end{cases}$
$k=5$.
$\begin{cases}v_{2}=X^{11},\cr v_{3}=X^{11}+X^{12},\cr v_{4}=X^{8}+X^{10}+X^{11},\cr v_{5}=X^{3}+X^{9}.\end{cases}$
$k=6$.
$\begin{cases}v_{2}=X^{11},\cr v_{3}=X^{11}+X^{12},\cr v_{4}=X^{10},\cr v_{5}=X^{6}+X^{7}+X^{8}+X^{9},\cr v_{6}=X+X^{2}+X^{3}+X^{4}+X^{8}.\end{cases}$
Note. To prove $6\in\mathcal{K}_{13}$, we could also use §2.2 Result (5): $\in\mathcal{K}_{13}, 33$$\cdot 3<13$ $\Rightarrow$ $3+3=6\in\mathcal{K}_{13}$.
Let $E$ be an $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q^{n}}$ and define
(3.1) $L_{E}(X)=\prod_{u\in E}(X-u)\in\mathbb{F}_{q^{n}}[X].$
It is well known ([16, Theorem 3.52]) that $L_{E}(X)$ is a $q$-polynomial, i.e., of the form $\sum_{i=0}^{d}a_{i}X^{q^{i}}$; hence $L_{E}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is an $\mathbb{F}_{q}$-linear map with kernel $E$.
Lemma 3.2.
Let $E=\mathbb{F}_{2^{l}}\subset\mathbb{F}_{2^{n}}$. Let $U\subset L_{E}(\mathbb{F}_{2^{n}})$ be a subspace with $\dim_{\mathbb{F}_{2}}U=r$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}U=s$, where $\mathbb{F}_{2^{l}}U$ is the $\mathbb{F}_{2^{l}}$-span of $U$.
  • (i)
    Let $V\subset\mathbb{F}_{2^{n}}$ be a subspace such that
    (3.2) $\dim_{\mathbb{F}_{2}}V=r\quad\text{and}\quad L_{E}(V)=U.$
    (Note that condition (3.2) implies $V\cap E=\{0\}$.) Then $s\leq\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V\leq s+1$.
  • (ii)
    We have $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=s$ for some $V$ satisfying condition (3.2).
Note that in Lemma 3.2, $s\leq r$, which gives Corollary 3.4 later.
Proof of Lemma 3.2.
(i) We have a short exact sequence of $\mathbb{F}_{2^{l}}$-subspaces
$0\longrightarrow(\mathbb{F}_{2^{l}}V)\cap E\longrightarrow\mathbb{F}_{2^{l}}V\xrightarrow{\ L_{E}\ }L_{E}(\mathbb{F}_{2^{l}}V)\longrightarrow 0,$
where $L_{E}(\mathbb{F}_{2^{l}}V)=\mathbb{F}_{2^{l}}L_{E}(V)=\mathbb{F}_{2^{l}}U$. Hence
(3.3) $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}U+\dim_{\mathbb{F}_{2^{l}}}(\mathbb{F}_{2^{l}}V)\cap E=\begin{cases}s&\text{if}\ E\not\subset\mathbb{F}_{2^{l}}V,\cr s+1&\text{if}\ E\subset\mathbb{F}_{2^{l}}V.\end{cases}$
(ii) Since $L_{E}:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is an $\mathbb{F}_{2^{l}}$-map and $\mathbb{F}_{2^{l}}U$ is an $\mathbb{F}_{2^{l}}$-subspace of $L_{E}(\mathbb{F}_{2^{n}})$, there exists an $\mathbb{F}_{2^{l}}$-subspace $W$ of $\mathbb{F}_{2^{n}}$ with $\dim_{\mathbb{F}_{2^{l}}}W=s$ such that $L_{E}(W)=\mathbb{F}_{2^{l}}U$. Note that $L_{E}|_{W}:W\to\mathbb{F}_{2^{l}}U$ is an $\mathbb{F}_{2^{l}}$-isomorphism. Let $V=(L_{E}|_{W})^{-1}(U)$. Then $L_{E}(V)=U$, $\dim_{\mathbb{F}_{2}}V=\dim_{\mathbb{F}_{2}}U=r$, and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V\leq\dim_{\mathbb{F}_{2^{l}}}W=s$. ∎
Remark. In Lemma 3.2 (i),
$\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=s+1\ \Leftrightarrow\ E\cap V=\{0\},\ E\subset\mathbb{F}_{2^{l}}V\ \text{and}\ U=L_{E}(V).$
Subspaces $V$ satisfying the above conditions do exist. For example, assume $\mathbb{F}_{2}\subsetneq E=\mathbb{F}_{2^{l}}\subsetneq\mathbb{F}_{2^{n}}$. Let $\alpha\in E\setminus\mathbb{F}_{2}$ and $\beta\in\mathbb{F}_{2^{n}}\setminus E$, and let $V$ be the $\mathbb{F}_{2}$-span of $\{\beta,\alpha+\alpha\beta\}$. If $a,b\in\mathbb{F}_{2}$ are such that $a\beta+b(\alpha+\alpha\beta)\in E\cap V$, then $(a+b\alpha)\beta+b\alpha\in E$, whence $a+b\alpha=0$, i.e., $a=b=0$. Therefore, $E\cap V=\{0\}$. Since $\alpha=(-\alpha)\beta+(\alpha+\alpha\beta)\in\mathbb{F}_{2^{l}}V$, we have $E=\mathbb{F}_{2^{l}}\alpha\subset\mathbb{F}_{2^{l}}V$.
Theorem 3.3.
Assume $l\geq 2$, $l\mid n$, and that there is an $r$-dimensional $\mathbb{F}_{2}$-subspace $F$ of $\mathbb{F}_{2^{n}}$ such that $\sum_{0\neq v\in F}1/v=0$. Let $s=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F$. Then for each $0\leq t\leq n/l-s$, there is a $(tl+r)$-dimensional $\mathbb{F}_{2}$-subspace $F_{t}$ of $\mathbb{F}_{2^{n}}$ such that $\sum_{0\neq v\in F_{t}}1/v=0$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s$. In particular,
$\Bigl\{tl+r:0\leq t\leq\frac{n}{l}-s\Bigr\}\subset\mathcal{K}_{n}.$
Proof.
Let $E=\mathbb{F}_{2^{l}}$. We use induction on $t$. For $t=0$, let $F_{0}=F$.
Assume that $F_{t}$ is given, where t<n/lst<n/l-s. Since $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s<n/l$, there exists $0\neq a\in\mathbb{F}_{2^{n}}$ such that $\text{\rm Tr}_{2^{n}/2^{l}}(av)=0$ for all $v\in\mathbb{F}_{2^{l}}F_{t}$, equivalently, $\text{\rm Tr}_{2^{n}/2^{l}}(av)=0$ for all $v\in F_{t}$. Thus $aF_{t}\subset L_{E}(\mathbb{F}_{2^{n}})$. Note that $\dim_{\mathbb{F}_{2}}aF_{t}=tl+r$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}(aF_{t})=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s$. By Lemma 3.2 (ii), there is an $\mathbb{F}_{2}$-subspace $V\subset\mathbb{F}_{2^{n}}$ such that $L_{E}(V)=aF_{t}$, $\dim_{\mathbb{F}_{2}}V=tl+r$ (hence $V\cap E=\{0\}$) and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=t+s$ (hence $(\mathbb{F}_{2^{l}}V)\cap E=\{0\}$ by equation (3.3)). Let $F_{t+1}=E\oplus V$. Then $\dim_{\mathbb{F}_{2}}F_{t+1}=l+tl+r=(t+1)l+r$,
$\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t+1}\,$ $=1+\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V\quad\text{(since $E\cap(\mathbb{F}_{2^{l}}V)=\{0\}$)}$
$=t+1+s.$
Moreover, by [5, Corollary 1], with $b=\prod_{0\neq u\in E}u$,
$\sum_{0\neq u\in F_{t+1}}\frac{1}{u}=\sum_{0\neq u\in E\oplus V}\frac{1}{u}=\sum_{0\neq v\in L_{E}(V)}\frac{b}{v}=\sum_{0\neq v\in F_{t}}\frac{b}{av}=\frac{b}{a}\sum_{0\neq v\in F_{t}}\frac{1}{v}=0.$
Corollary 3.4.
Let $l\geq 2$, $l\mid n$. If $r\in\mathcal{K}_{n}$, then
$\Bigl\{tl+r:0\leq t\leq\frac{n}{l}-r\Bigr\}\subset\mathcal{K}_{n}.$
Proof.
Note that in Theorem 3.3, $s\leq r$. ∎
Example 3.5.
Let $f=X^{21}+X^{19}+1\in\mathbb{F}_{2}[X]$, which is irreducible (in fact, primitive) over $\mathbb{F}_{2}$. Hence $\mathbb{F}_{2^{21}}=\mathbb{F}_{2}[X]/(f)$. We search for $1,v_{2},v_{3},v_{4}\in\mathbb{F}_{2^{21}}$ with $v_{2}\in\mathbb{F}_{2^{3}}$ such that $\Delta(1,v_{2},v_{3},v_{4})\neq 0$ but $\Delta_{1}(1,v_{2},v_{3},v_{4})=0$. The following is an example:
$\begin{cases}v_{2}=X^{19}+X^{16}+X^{15}+X^{14}+X^{11}+X^{9}+X^{7}+X^{6}+X^{4}+X^{3}+X^{2}+1,\cr v_{3}=X^{20}+X^{19}+X^{18}+X^{13}+X^{11}+X^{10}+X^{9}+X^{8}+X^{4}+X^{3}+1,\cr v_{4}=X^{20}+X^{19}+X^{16}+X^{15}+X^{14}+X^{13}+X^{12}+X^{11}+X^{10}+X^{8}+X^{4}+X^{3}+1.\end{cases}$
Let $F$ be the $\mathbb{F}_{2}$-subspace with basis $\{1,v_{2},v_{3},v_{4}\}$. Then $\dim_{\mathbb{F}_{2^{3}}}\mathbb{F}_{2^{3}}F\leq 3$ since $v_{2}\in\mathbb{F}_{2^{3}}$. Now by Theorem 3.3 with $n=21$, $l=3$, $r=4$, $s\leq 3$ and $t=4\leq n/l-s$, we have $tl+r=4\cdot 3+4=16\in\mathcal{K}_{21}$. Hence $5=21-16\in\mathcal{K}_{21}$.
Note. In fact, Conjecture 1.1 is true for $n=21$ by [7, Table 2].
Theorem 3.6.
Let $n\geq 3$ be odd and $l$ be the smallest prime divisor of $n$. If
(3.4) $(l-1)(l+2)\leq\frac{n+1}{2},$
then Conjecture 1.1 is true for $n$.
Note. Inequality (3.4) is equivalent to $n\geq 2l(l+1)-5$. When $l>5$, this is equivalent to $n\geq 2l(l+1)$ since $l\mid n$.
Proof of Theorem 3.6.
We may assume $n>13$ since the conjecture is known to be true for $n\leq 13$.
$1^{\circ}$ First assume
$l+2\leq\frac{3n-9}{13}.$
By [7, Theorem 4.2], $3,4,\dots,l+2\in\mathcal{K}_{n}$. Then by Theorem 3.3, for $3\leq r\leq l+2$,
$tl+r\in\mathcal{K}_{n}\quad\text{for}\quad 0\leq t\leq\frac{n}{l}-r.$
Since
$\Bigl(\frac{n}{l}-r\Bigr)l+r=n-(l-1)r\geq n-(l-1)(l+2)\geq n-\frac{n+1}{2}=\frac{n-1}{2},$
we see that
$\mathcal{K}_{n}\supset\Bigl\{tl+r:3\leq r\leq l+2,\ 0\leq t\leq\frac{n}{l}-r\Bigr\}\supset\Bigl\{3,4,\dots,\frac{n-1}{2}\Bigr\}.$
Since $k\in\mathcal{K}_{n}$ if and only if $n-k\in\mathcal{K}_{n}$, the conjecture holds in this case.
$2^{\circ}$ Now assume
(3.5) $l+2>\frac{3n-9}{13}.$
Then
$\Bigl(\frac{3n-9}{13}-3\Bigr)\frac{3n-9}{13}<\frac{n+1}{2},$
which implies that $n<27$. The pair $(n,l)$ satisfying inequalities (3.4), where $n$ is odd and $13<n<27$, is $(21,3)$. By [7, Table 2], Conjecture 1.1 is true for $n=21$. ∎
Corollary 3.7.
If $3\mid n$, then Conjecture 1.1 is true for $n$.
Proof.
We may assume that $n>13$ since the conjecture is true for $n\leq 13$. By Theorem 3.6, we may assume that
$(3-1)(3+2)>\frac{n+1}{2},$
i.e., $n<19$. Since $3\mid n$, we then have $n=15$. By [7, Table 2], Conjecture 1.1 is true for $n=15$. ∎
In [5], Carlet established the following criterion for $k\in\mathcal{K}_{n}$ using the indicator function of an (nk)(n-k)-dimensional subspace of $\mathbb{F}_{2^{n}}$:
Theorem 3.8 ([5, §5.2.4]).
We have $k\in\mathcal{K}_{n}$ if and only if there exist $u_{1},\dots,u_{k}\in\mathbb{F}_{2^{n}}$, linearly independent over $\mathbb{F}_{2}$, such that
$\theta(u_{1},\dots,u_{k})=0,$
where
(3.6) $\theta(u_{1},\dots,u_{k})=\sum_{\begin{subarray}{c}b_{1},\dots,b_{k}\in\{-\infty,0,\dots,n-1\}\cr 2^{b_{1}}+\cdots+2^{b_{k}}\equiv 1\,(\text{\rm mod}\;2^{n}-1)\end{subarray}}u_{1}^{2^{b_{1}}}\cdots u_{k}^{2^{b_{k}}},$
and $2^{-\infty}$ is defined to be 0.
This criterion can be reformulated in terms of symmetric polynomials.
Lemma 3.9.
Assume that $2\leq l\leq 2^{n}$. Then
$\text{\rm wt}_{2}(1+l(2^{n}-1))=n+1-\nu_{2}(l-1),$
where $\text{\rm wt}_{2}$ is the base $2$ weight and $\nu_{2}$ is the $2$-adic order.
Proof.
Write
$1+l(2^{n}-1)=2^{n}-(l-1)+2^{n}(l-1).$
Let the base 2 digits of l1l-1 be
$\begin{matrix}\scriptstyle 0&\scriptstyle 1&\cdots&\scriptstyle i-1&\scriptstyle i&\scriptstyle i+1&\cdots&\scriptstyle n-1\cr 0&0&\cdots&0&1&a_{i+1}&\cdots&a_{n-1}\end{matrix}\ ,$
where $i=\nu_{2}(l-1)$. Then the base 2 digits of $2^{n}-(l-1)$ are
$\begin{matrix}\scriptstyle 0&\scriptstyle 1&\cdots&\scriptstyle i-1&\scriptstyle i&\scriptstyle i+1&\cdots&\scriptstyle n-1\cr 0&0&\cdots&0&1&\overline{a_{i+1}}&\cdots&\overline{a_{n-1}}\end{matrix}\ ,$
where $\overline{a_{j}}=1-a_{j}$. Therefore,
$\text{wt}_{2}(1+l(2^{n}-1))=\text{wt}_{2}(2^{n}-(l-1))+\text{wt}_{2}(l-1)=2+n-1-i=n+1-i.$
Lemma 3.10.
Assume $k\leq n$ and $b_{1},\dots,b_{k}\in\{0,1,\dots,n-1\}$ are such that
$2^{b_{1}}+\cdots+2^{b_{k}}\equiv 1\pmod{2^{n}-1}$.
Then $2^{b_{1}}+\cdots+2^{b_{k}}=1$ or $2^{n}$.
Proof.
Assume to the contrary that
(3.7) $2^{b_{1}}+\cdots+2^{b_{k}}=1+l(2^{n}-1)$
for some $l\geq 2$. Then $1+l(2^{n}-1)\leq k2^{n-1}\leq n2^{n-1}$ and hence
$l\leq\frac{n2^{n-1}-1}{2^{n}-1}\leq 2^{n}.$
By Lemma 3.9,
(3.8) $\text{wt}_{2}(1+l(2^{n}-1))=n+1-\nu_{2}(l-1).$
Since
$2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}-(l-1)+2^{n}(l-1),$
the sum at the left hand side has at least l1l-1 carries. Consequently,
(3.9) $\text{wt}_{2}(2^{b_{1}}+\cdots+2^{b_{k}})\leq k-(l-1).$
Combining equations (3.7), (3.8) and inequality (3.9), we have
$k-(l-1)\geq n+1-\nu_{2}(l-1)\geq n+1-\log_{2}(l-1).$
Hence
$k\geq n+1+l-1-\log_{2}(l-1)\geq n+1,$
which is a contradiction. ∎
By Lemma 3.10, equation (3.6) becomes
(3.10) $\theta(u_{1},\dots,u_{k})=u_{1}+\cdots+u_{k}+\sum_{\begin{subarray}{c}b_{1},\dots,b_{k}\in\{-\infty,0,\dots,n-1\}\cr 2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}\end{subarray}}u_{1}^{2^{b_{1}}}\cdots u_{k}^{2^{b_{k}}}.$
Assume $k\geq 2$ and $2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}$. Then $\min\{b_{i}:b_{i}\neq-\infty\}\geq n-k+1$ and the equality holds if and only if the multiset $\{b_{1},\dots,b_{k}\}$ equals $\{n-k+1,n-k+1,n-k+2,\dots,n-1\}$. Thus by equation (3.10),
$\theta(u_{1},\dots,u_{k})^{2^{k-1}}=u_{1}^{2^{k-1}}+\cdots+u_{k}^{2^{k-1}}+\sum_{\begin{subarray}{c}b_{1},\dots,b_{k}\in\{-\infty,n-k+1,\dots,n-1\}\cr 2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}\end{subarray}}u_{1}^{2^{b_{1}^{\prime}}}\cdots u_{k}^{2^{b_{k}^{\prime}}},$
where $b_{i}^{\prime}=b_{i}-(n-k+1)$. ($\infty-(n-k+1)=-\infty$.) Note that $b_{1},\dots,b_{k}\in\{-\infty,n-k+1,\dots,n-1\}$ satisfy
$2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}$
if and only if $b_{1}^{\prime},\dots,b_{k}^{\prime}\in\{-\infty,0,\dots,k-2\}$ satisfy
$2^{b_{1}^{\prime}}+\cdots+2^{b_{k}^{\prime}}=2^{-(n-k+1)}2^{n}=2^{k-1}.$
Therefore,
$\theta(u_{1},\dots,u_{k})^{2^{k-1}}=\sum_{l=1}^{k}\,\sum_{\begin{subarray}{c}2^{c_{1}}\geq\cdots\geq 2^{c_{l}}\cr 2^{c_{1}}+\cdots+2^{c_{l}}=2^{k-1}\end{subarray}}\,\sum_{\begin{subarray}{c}(i_{1},\dots,i_{l})\cr i_{1},\dots,i_{l}\in\{1,\dots,k\}\;\text{distinct}\end{subarray}}u_{i_{1}}^{2^{c_{1}}}\cdots u_{i_{l}}^{2^{c_{l}}}.$
In the above,
$\sum_{\begin{subarray}{c}(i_{1},\dots,i_{l})\cr i_{1},\dots,i_{l}\in\{1,\dots,k\}\;\text{distinct}\end{subarray}}u_{i_{1}}^{2^{c_{1}}}\cdots u_{i_{l}}^{2^{c_{l}}}=m_{(2^{c_{1}},\dots,2^{c_{l}})}(u_{1},\dots,u_{k}).$
where $m_{(2^{c_{1}},\dots,2^{c_{l}})}(X_{1},\dots,X_{k})$ is the monomial symmetric polynomial associated to the partition $(2^{c_{1}},\dots,2^{c_{l}})$ ([17, I.2]). We shall call a partition (of a non-negative integer) a $2$-adic partition if the parts are powers of $2$. Thus,
$\theta(u_{1},\dots,u_{k})^{2^{k-1}}=\sum_{\lambda}m_{\lambda}(u_{1},\dots,u_{k}),$
where $\lambda$ runs through all $2$-adic partitions of $2^{k-1}$ with at most $k$ parts, i.e., $\lambda=(2^{c_{1}},\dots,2^{c_{l}})\vdash 2^{k-1}$, $l\leq k$. Let $\Lambda_{k}$ denote the set of all $2$-adic partitions of $2^{k-1}$ with at most $k$ parts and define
$\Theta_{k}(X_{1},\dots,X_{k})=\sum_{\lambda\in\Lambda_{k}}m_{\lambda}(X_{1},\dots,X_{k}).$
With this notation, Theorem 3.8 is reformulated as follows:
Theorem 3.11.
We have $k\in\mathcal{K}_{n}$ if and only if there are $u_{1},\dots,u_{k}\in\mathbb{F}_{2^{n}}$ such that $\Delta(u_{1},\dots,u_{k})\neq 0$ but $\Theta_{k}(u_{1},\dots,u_{k})=0$.
Elements of $\Lambda_{k}$ are of the form $\lambda=(\underbrace{2^{k-1},\dots,2^{k-1}}_{a_{k-1}},\dots,\underbrace{2^{0},\dots,2^{0}}_{a_{0}})$, where
$\begin{cases}a_{0}2^{0}+\cdots+a_{k-1}2^{k-1}=2^{k-1},\cr a_{0}+\cdots+a_{k-1}\leq k.\end{cases}$
For example, for $\Lambda_{4}$,
$(a_{0},a_{1},a_{2},a_{3})=\,$ $(0,0,0,1),$
$(0,0,2,0),$
$(0,2,1,0),$
$(2,1,1,0),$
$(0,4,0,0).$
So,
$\Theta_{4}=m_{(2^{3})}+m_{(2^{2},2^{2})}+m_{(2^{2},2,2)}+m_{(2^{2},2,1,1)}+m_{(2,2,2,2)}.$
The enumeration of all terms in $\Theta_{4}$ are given in the appendix.
The next theorem is an application of Theorem 3.11
Theorem 3.12.
For $n\geq 6$, $4\in\mathcal{K}_{n}$.
Proof.
For $n\leq 15$, the claim is known to be true. (The cases $n=13,15$ follow from Theorem 3.1 and Corollary 3.7.) Therefore, we assume $n\geq 17$.
Let $q=2^{n}$ and for $f\in\mathbb{F}_{q}[X_{1},\dots,X_{4}]$, let
$V_{\mathbb{F}_{q}^{4}}(f)=\{(x_{1},\dots,x_{4})\in\mathbb{F}_{q}^{4}:f(x_{1},\dots,x_{4})=0\}.$
By Theorem 3.11, it suffices to show that $V_{\mathbb{F}_{q}^{4}}(\Theta_{4})\not\subset V_{\mathbb{F}_{q}^{4}}(\Delta)$. We are able to show that $\Theta_{4}(X_{1},\dots,X_{4})$ is irreducible over $\overline{\mathbb{F}}_{2}$; see Theorem A1 in the appendix. Thus by the Lang-Weil bound [2, Theorem 5.2],
$|V_{\mathbb{F}_{q}^{4}}(\Theta_{4})|\geq q^{3}-(8-1)(8-2)q^{5/2}-5\cdot 8^{13/3}q^{2}=q^{3}-42q^{5/2}-5\cdot 2^{13}q^{2},$
where $8=\deg\Theta_{4}$. By [2, Lemma 2.2],
$|V_{\mathbb{F}_{q}^{4}}(\Theta_{4})\cap V_{\mathbb{F}_{q}^{4}}(\Delta)|\leq 15^{2}q^{2},$
where $15=\deg\Delta$. Hence
$|V_{\mathbb{F}_{q}^{4}}(\Theta_{4})\setminus V_{\mathbb{F}_{q}^{4}}(\Delta)|\geq q^{3}-42q^{5/2}-5\cdot 2^{13}q^{2}-15q^{2}=q^{2}(q-42q^{1/2}-41185).$
Therefore, it suffices to show that
(3.11) $q-42q^{1/2}-41185>0.$
The larger root of $X^{2}-42X-41185$ is
$\frac{1}{2}(42+\sqrt{42^{2}+4\cdot 41185})=21+\sqrt{41626}.$
Thus, inequality (3.11) is satisfied when
$q^{1/2}=2^{n/2}>21+\sqrt{41626},$
i.e.,
$n>2\log_{2}(21+\sqrt{41626})\approx 15.63.$
This completes the proof. ∎
Theorem 3.13.
If $5\mid n$, then Conjecture 1.1 is true for $n$.
Proof.
We may assume that $n\geq 17$. By Theorem 3.6, we may assume that (51)(5+2)>(n+1)/2(5-1)(5+2)>(n+1)/2, i.e., $n<55$. By Corollary 3.7, we may further assume $3\nmid n$. Therefore, we only have to consider $n=25,35$.
When $n=25$, we only have to check that $3,\dots,12\in\mathcal{K}_{25}$. We already know that $3,4,5\in\mathcal{K}_{25}$ (§2.2 Results (1), (6), and Theorem 3.12). Using §2.2 Result (5) repeatedly, we see that $6=3+3,\,7=3+4,\,\dots,\,11=3+8$ all belong to $\mathcal{K}_{25}$. By Corollary 3.4, $13=2\cdot 5+3\in\mathcal{K}_{25}$, whence $12=25-13\in\mathcal{K}_{25}$.
When $n=35$, we only have to check that $3,\dots,17\in\mathcal{K}_{35}$. As in the above, we have $3,\,4,\,5,\,6(=3+3),\,\dots,\,14(=3+11)\in\mathcal{K}_{35}$. Also, $15\in\mathcal{K}_{35}$ since $\gcd(15,35)>1$. By Corollary 3.4, both $18=3\cdot 5+3$ and $19=3\cdot 5+4$ belong to $\mathcal{K}_{35}$. Therefore, $16=35-19\in\mathcal{K}_{35}$ and $17=35-18\in\mathcal{K}_{35}$. ∎
4. Generalizations
Recall that a function $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional $\mathbb{F}_{q}$-affine subspaces $A$ of $\mathbb{F}_{q^{n}}$. In general, compared with their binary counterpart, $q$-ary sum-free functions are more difficult to characterize, construct and analyze. APN functions in odd characteristic, studied in [9, 10], do not represent the 2nd order $q$-ary sum-free functions. On the other hand, the generalized almost perfect nonlinear (GAPN) functions in odd characteristic, introduced and explored in [15], appear to be related to the 1st order sum-free functions. However, this connection will not be pursued in the present paper.
We first take a look of the 1st order and the (n1)(n-1)th order $q$-ary sum-free functions.
4.1. The 1st order
Let $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ be a function represented by $f(X)=\sum_{i=0}^{q^{n}-1}a_{i}X^{i}\in\mathbb{F}_{q^{n}}[X]$. Then $f$ is 1st order sum-free if and only if $\sum_{x\in\mathbb{F}_{q}}f(b(x+c))\neq 0$ for all $b\in\mathbb{F}_{q^{n}}^{*}$ and $c\in\mathbb{F}_{q^{n}}$. We have
$\sum_{x\in\mathbb{F}_{q}}f(b(x+c))=\sum_{i=0}^{q^{n}-1}a_{i}b^{i}\sum_{x\in\mathbb{F}_{q}}(x+c)^{i}.$
There is a polynomial $g_{i,q}\in\mathbb{F}_{p}[X]$ ($p=\text{char}\,\mathbb{F}_{q}$) such that
(4.1) $\sum_{x\in\mathbb{F}_{q}}(x+X)^{i}=g_{i,q}(X^{q}-X).$
The polynomial $g_{i,q}$ was introduced and studied in [12] for a different purpose and it satisfies the recursive relation
$\begin{cases}g_{0,q}=\cdots=g_{q-2,q}=0,\cr g_{q-1,q}=-1,\cr g_{i,q}=Xg_{i-q,q}+g_{i-q+1,q},\quad i\geq q.\end{cases}$
Therefore,
$\sum_{x\in\mathbb{F}_{q}}f(b(x+c))=\sum_{i=0}^{q^{n}-1}a_{i}b^{i}g_{i,q}(c^{q}-c).$
Consequently, we have the following proposition.
Proposition 4.1.
Let $f(X)=\sum_{i=0}^{q^{n}-1}a_{i}X^{i}\in\mathbb{F}_{q^{n}}[X]$. Then $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is 11st order sum-free if and only if
$\sum_{i=0}^{q^{n}-1}a_{i}b^{i}g_{i,q}(d)\neq 0$
for all $b\in\mathbb{F}_{q^{n}}^{*}$ and $d\in\mathbb{F}_{q^{n}}$ with $\text{\rm Tr}_{q^{n}/q}(d)=0$.
In particular, a power function $X^{i}$ is 1st order sum-free on $\mathbb{F}_{q^{n}}$ if and only if $g_{i,q}(d)\neq 0$ for all $d\in\mathbb{F}_{q^{n}}$ with $\text{\rm Tr}_{q^{n}/q}(d)=0$. Here, when $d=c^{q}-c$ ($c\in\mathbb{F}_{q^{n}}$),
$g_{i,q}(d)=\sum_{x\in\mathbb{F}_{q}}(x+c)^{i}=-\sum_{0<j\equiv 0\,\text{(mod $q-1$)}}\binom{i}{j}c^{i-j}.$
Example 4.2.
Let $q=3$ and $n=2$. We have
$g_{i,3}(X)=\begin{cases}0&\text{if}\ i=0,1,3,\cr-1&\text{if}\ i=2,4,6,\cr-X&\text{if}\ i=5,\cr X&\text{if}\ i=7,\cr-X^{2}-1&\text{if}\ i=8.\end{cases}$
For $0\leq i\leq 8$, $\gcd(g_{i,3}(X^{3}-X),X^{9}-X)=1$ if and only if $i=2,4,6$. Hence $X^{i}$ is 1st order sum-free on $\mathbb{F}_{3^{2}}$ if and only if $i\in\{2,4,6\}$.
4.2. The (n1)(n-1)th order
Let $f(X)=\sum_{i=0}^{q^{n}-1}a_{i}X^{i}\in\mathbb{F}_{q^{n}}[X]$, where $n\geq 2$, and let $b\in\mathbb{F}_{q^{n}}^{*}$ and $c\in\mathbb{F}_{q}$. The following computation generalizes that in the proof of Theorem 2.1. We have
$\sum_{\text{\rm Tr}_{q^{n}/q}(bx)=c}f(x)$
$=\sum_{x\in\mathbb{F}_{q^{n}}}f(x)\bigl[1-\bigl(\text{\rm Tr}_{q^{n}/q}(bx)-c\bigr)^{q-1}\bigr]$
$=\sum_{x\in\mathbb{F}_{q^{n}}}f(x)-\sum_{x\in\mathbb{F}_{q^{n}}}f(x)\sum_{j=0}^{q-1}\text{\rm Tr}_{q^{n}/q}(bx)^{j}c^{q-1-j}$
$=-a_{q^{n}-1}-\sum_{x\in\mathbb{F}_{q^{n}}}f(x)\sum_{j=0}^{q-1}c^{q-1-j}\bigl((bx)^{q^{0}}+\cdots+(bx)^{q^{n-1}}\bigr)^{j}$
$=-a_{q^{n}-1}-\sum_{\begin{subarray}{c}x\in\mathbb{F}_{q^{n}}\cr 0\leq j\leq q-1\end{subarray}}f(x)c^{q-1-j}\sum_{j_{0}+\cdots+j_{n-1}=j}\binom{j_{0}+\cdots+j_{n-1}}{j_{0},\dots,j_{n-1}}(bx)^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
$=-a_{q^{n}-1}-\sum_{j_{0}+\cdots+j_{n-1}\leq q-1}\binom{j_{0}+\cdots+j_{n-1}}{j_{0},\dots,j_{n-1}}c^{q-1-(j_{0}+\cdots+j_{n-1})}b^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
$\kern 85.35826pt\cdot\Bigl[\sum_{i=0}^{q^{n}-1}a_{i}\sum_{x\in\mathbb{F}_{q^{n}}}x^{i+j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}\Bigr]$
$=-a_{q^{n}-1}+\sum_{j_{0}+\cdots+j_{n-1}\leq q-1}\binom{j_{0}+\cdots+j_{n-1}}{j_{0},\dots,j_{n-1}}a_{q^{n}-1-(j_{0}q^{0}+\cdots+j_{n-1}q^{n-1})}b^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
$\kern 85.35826pt\cdot c^{q-1-(j_{0}+\cdots+j_{n-1})}.$
When $c=0$,
(4.2) $\sum_{\text{\rm Tr}_{q^{n}/q}(bx)=0}f(x)=$
$-a_{q^{n}-1}+\sum_{j_{0}+\cdots+j_{n-1}=q-1}\binom{q-1}{j_{0},\dots,j_{n-1}}a_{q^{n}-1-(j_{0}q^{0}+\cdots+j_{n-1}q^{n-1})}b^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}.$
When $c\neq 0$,
(4.3) $\sum_{\text{\rm Tr}_{q^{n}/q}(bx)=c}f(x)=$
(4.4) $\sum_{0<j_{0}+\cdots+j_{n-1}\leq q-1}\binom{j_{0}+\cdots+j_{n-1}}{j_{0},\dots,j_{n-1}}a_{q^{n}-1-(j_{0}q^{0}+\cdots+j_{n-1}q^{n-1})}b^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
$\kern 85.35826pt\cdot c^{q-1-(j_{0}+\cdots+j_{n-1})}.$
Therefore, we have the following proposition.
Proposition 4.3.
In the above notation, $f$ is (n1)(n-1)th order sum-free if and only if
(4.5) $a_{q^{n}-1}\neq\sum_{j_{0}+\cdots+j_{n-1}=q-1}\binom{q-1}{j_{0},\dots,j_{n-1}}a_{q^{n}-1-(j_{0}q^{0}+\cdots+j_{n-1}q^{n-1})}b^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
for all $b\in\mathbb{F}_{q^{n}}^{*}$, and
(4.6) $\sum_{0<j_{0}+\cdots+j_{n-1}\leq q-1}\binom{j_{0}+\cdots+j_{n-1}}{j_{0},\dots,j_{n-1}}a_{q^{n}-1-(j_{0}q^{0}+\cdots+j_{n-1}q^{n-1})}b^{j_{0}q^{0}+\cdots+j_{n-1}q^{n-1}}$
$\kern 85.35826pt\cdot c^{q-1-(j_{0}+\cdots+j_{n-1})}\neq 0$
for all $b\in\mathbb{F}_{q^{n}}^{*}$ and $c\in\mathbb{F}_{q}^{*}$.
Condition (4.5) can always be satisfied since the right side of inequality (4.5) takes at most $q^{n}-1$ values as $b$ runs over $\mathbb{F}_{q^{n}}^{*}$. Condition (4.6) is difficult to verify in general. However, in special cases, condition (4.6) can be made more explicit.
Let $0<t\leq q-1$ and assume that for $0<j_{0}+\cdots+j_{n-1}\leq q-1$,
$a_{q^{n}-1-(j_{0}q^{0}+\cdots+j_{n-1}q^{n-1})}=0$
unless
[…]
where $p=\text{char}\,\mathbb{F}_{q}$, or
$(j_{0},\dots,j_{n-1})=(0,\dots,0,\overset{i}{t},0\,\dots,0)\quad\text{for some }\ 0\leq i\leq n-1.$
Then the left side of inequality (4.6) becomes
$c^{q-1-t}\sum_{i=0}^{n-1}a_{q^{n}-1-tq^{i}}b^{tq^{i}}=c^{q-1-t}L_{t}(b^{t}),$
where
$L_{t}(X)=\sum_{i=0}^{n-1}a_{q^{n}-1-tq^{i}}X^{q^{i}}$
is a $q$-polynomial. Therefore, if $L_{t}(X)$ is a permutation polynomial of $\mathbb{F}_{q^{n}}$, then condition (4.6) is satisfied.
Example 4.4.
Let $q=5$, $n=2$, and $\mathbb{F}_{5^{2}}=\mathbb{F}_{5}(\alpha)$, where $\alpha^{2}=2$. Choose a permutation $q$-polynomial of $\mathbb{F}_{q^{2}}$, say $X+\alpha X^{5}$. Let $t=3$ and choose $a_{24-3}$ and $a_{24-3\cdot 5}$ such that
$X+\alpha X^{5}=L_{3}(X)=a_{24-3}X+a_{24-3\cdot 5}X^{5},$
i.e., $a_{21}=1$ and $a_{9}=\alpha$. For $i\in\{0,\dots,23\}\setminus\{9,21\}$, set $a_{i}=0$. Then codition (4.6) is satisfied. Moreover, the right side of (4.5) is 0. Therefore, for all $a_{24}\in\mathbb{F}_{5^{2}}^{*}$,
$f(X)=\alpha X^{9}+X^{21}+a_{24}X^{24}$
is 1st order sum-free on $\mathbb{F}_{5^{2}}$.
For the rest of Section 4, we focus on the $q$-ary version of the binary multiplicative inverse function $f_{\text{\rm inv}}$. First, we have to determine the “right” $q$-ary generalization of $f_{\text{\rm inv}}$. To this end, we need to recall some formulas related to Moore determinant.
4.3. Formulas about the Moore determinant
Recall that the $q$-ary Moore determinant $\Delta(X_{1},\dots,X_{k})$ and its relatives $\Delta_{i}(X_{1},$ $\dots,X_{k})$ are defined in §2.3.
Lemma 4.5.
  • (i)
    We have
    (4.7) $\prod_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})=(-1)^{k}\frac{\Delta(Y,X_{1},\dots,X_{k})}{\Delta(X_{1},\dots,X_{k})}.$
  • (ii)
    Let $\Delta=\Delta(Y,X_{1},\dots,X_{k})$ and $\Delta_{i}=\Delta_{i}(X_{1},\dots,X_{k})$. For $l\geq 0$, we have
    (4.8) $\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{1}{(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})^{l+1}}$
    $=\frac{\Delta_{0}}{\Delta}\sum_{l_{0}q^{0}+\cdots+l_{k}q^{k}=l}\binom{l_{0}+\cdots+l_{k}}{l_{0},\dots,l_{k}}\prod_{i=0}^{k}\Bigl(\frac{(-1)^{i}\Delta_{i})}{\Delta}\Bigr)^{l_{i}}.$
Proof.
(i) $\Delta(Y,X_{1},\dots,X_{k})$ is a polynomial of degree $q^{k}$ in $Y$ with roots $(a_{1}X_{1}+\cdots+a_{k}X_{k})$, $a_{1},\dots,a_{k}\in\mathbb{F}_{q}$; $(-1)^{k}\Delta(X_{1},\dots,X_{k})$ is the leading coefficient.
(ii) Differentiating both sides of equation (4.7) with respect to $Y$ gives
(4.9) $\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{1}{Y+a_{1}X_{1}+\cdots+a_{k}X_{k}}=\frac{\Delta(X_{1},\dots,X_{k})^{q}}{\Delta(Y,X_{1},\dots,X_{k})}=\frac{\Delta_{0}}{\Delta}.$
Consider the generating function
$G_{k}(Z)=\sum_{l=1}^{\infty}\Bigl(\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{1}{(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})^{l}}\Bigr)Z^{l}.$
We have
$G_{k}(Z)\,$ $=\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\sum_{l=1}^{\infty}\Bigl(\frac{Z}{Y+a_{1}X_{1}+\cdots+a_{k}X_{k}}\Bigr)^{l}$
$=\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{\displaystyle\frac{Z}{Y+a_{1}X_{1}+\cdots+a_{k}X_{k}}}{\displaystyle 1-\frac{Z}{Y+a_{1}X_{1}+\cdots+a_{k}X_{k}}}$
$=Z\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{1}{Y-Z+a_{1}X_{1}+\cdots+a_{k}X_{k}}$
[…]
$=Z\,\frac{\Delta_{0}}{\Delta-\Delta(Z,X_{1},\dots,X_{k})}$
$=Z\,\frac{\Delta_{0}}{\Delta}\cdot\frac{1}{\displaystyle 1-\frac{\Delta(Z,X_{1},\dots,X_{k})}{\Delta}}$
$=Z\,\frac{\Delta_{0}}{\Delta}\sum_{l=0}^{\infty}\frac{\Delta(Z,X_{1},\dots,X_{k})^{l}}{\Delta^{l}}.$
In the above,
$\Delta(Z,X_{1},\dots,X_{k})^{l}=\Bigl(\sum_{i=0}^{k}(-1)^{i}\Delta_{i}Z^{q^{i}}\Bigr)^{l}$
$=\sum_{l_{0}+\cdots+l_{k}=l}\binom{l}{l_{0},\dots,l_{k}}\Bigl(\prod_{i=0}^{k}((-1)^{i}\Delta_{i})^{l_{i}}\Bigr)Z^{l_{0}q^{0}+\cdots+l_{k}q^{k}}.$
Therefore
$G_{k}(Z)=Z\,\frac{\Delta_{0}}{\Delta}\sum_{l=0}^{\infty}\;\sum_{l_{0}q^{0}+\cdots+l_{k}q^{k}=l}\binom{l_{0}+\cdots+l_{k}}{l_{0},\dots,l_{k}}\Bigl(\prod_{i=0}^{k}\Bigl(\frac{(-1)^{i}\Delta_{i})}{\Delta}\Bigr)^{l_{i}}\Bigr)Z^{l}.$
Hence
$\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{1}{(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})^{l+1}}$
$=\frac{\Delta_{0}}{\Delta}\sum_{l_{0}q^{0}+\cdots+l_{k}q^{k}=l}\binom{l_{0}+\cdots+l_{k}}{l_{0},\dots,l_{k}}\prod_{i=0}^{k}\Bigl(\frac{(-1)^{i}\Delta_{i})}{\Delta}\Bigr)^{l_{i}}.$
Remark. Equation (4.8) leads to a formula by Carlitz on the power sum of reciprocals of polynomials over finite fields [8, Theorem 9.2].
4.4. What is the natural $q$-ary generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$?
For $1\leq s\leq q-1$, let $g_{s}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ be defined by
$g_{s}(x)=\begin{cases}1/x^{s}&\text{if}\ x\neq 0,\cr 0&\text{if}\ x=0.\end{cases}$
(If $q^{n}>s+1$, then $g_{s}(x)=x^{q^{n}-1-s}$ for all $x\in\mathbb{F}_{q^{n}}$.) We remind the reader that $g_{s}$ here is different from the polynomial $g_{i,q}$ in equation (4.1). By equation (4.8),
(4.10) $\sum_{a_{1},\dots,a_{k}\in\mathbb{F}_{q}}\frac{1}{(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})^{s}}=\Bigl(\frac{\Delta_{0}}{\Delta}\Bigr)^{s}.$
Hence, if $E$ is a $k$-dimensional $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q^{n}}$ and $y\in\mathbb{F}_{q^{n}}\setminus E$, then
$\sum_{u\in E}\frac{1}{(y+u)^{s}}\neq 0.$
We have
$\sum_{0\neq(a_{1},\dots,a_{k})\in\mathbb{F}_{q}^{k}}\frac{1}{(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})^{s}}=\Bigl(\frac{\Delta_{0}}{\Delta}\Bigr)^{s}-\frac{1}{Y^{s}}.$
Note that
$\Delta=\sum_{i=0}^{k}(-1)^{i}Y^{q^{i}}\Delta_{i}=Y\Delta_{0}\Bigl(1-Y^{q-1}\frac{\Delta_{1}}{\Delta_{0}}+\text{terms of higher degree in $Y$}\Bigr).$
Hence,
$\Delta^{-s}=Y^{-s}\Delta_{0}^{-s}\Bigl(1+sY^{q-1}\frac{\Delta_{1}}{\Delta_{0}}+\cdots\Bigr).$
Therefore,
$\sum_{0\neq(a_{1},\dots,a_{k})\in\mathbb{F}_{q}^{k}}\frac{1}{(Y+a_{1}X_{1}+\cdots+a_{k}X_{k})^{s}}\,$ $=Y^{-s}\Bigl(1+sY^{q-1}\frac{\Delta_{1}}{\Delta_{0}}+\cdots\Bigr)-\frac{1}{Y^{s}}$
$\equiv sY^{q-1-s}\frac{\Delta_{1}}{\Delta_{0}}\pmod{Y^{q-s}}$.
Setting $Y=0$ gives
$\sum_{0\neq(a_{1},\dots,a_{k})\in\mathbb{F}_{q}^{k}}\frac{1}{(a_{1}X_{1}+\cdots+a_{k}X_{k})^{s}}=\begin{cases}-\displaystyle\frac{\Delta_{1}}{\Delta_{0}}&\text{if}\ s=q-1,\vskip 1.99997pt\cr 0&\text{if}\ 1\leq s<q-1.\end{cases}$
Let $E$ be a $k$-dimensional $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q^{n}}$ with basis $v_{1},\dots,v_{k}$. Then
$\sum_{0\neq u\in E}\frac{1}{u^{s}}=\begin{cases}-\displaystyle\frac{\Delta_{1}(v_{1},\dots,v_{k})}{\Delta(v_{1},\dots,v_{k})^{q}}&\text{if}\ s=q-1,\vskip 1.99997pt\cr 0&\text{if}\ 1\leq s<q-1.\end{cases}$
When s<q1s<q-1, $g_{s}$ is never $k$th order sum-free. Therefore, the “right” $q$-ary generalization of $f_{\text{\rm inv}}$ is $g_{q-1}$. From the above, we know that $g_{q-1}$ is not $k$th order sum-free if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ such that $\Delta(v_{1},\dots,v_{k})\neq 0$ but $\Delta_{1}(v_{1},\dots,v_{k})=0$. We state this fact as a theorem since it is the foundation of some subsequent results.
Theorem 4.6 ($q$-ary version of Theorem 2.4).
The function $g_{q-1}$ is not $k$th order sum-free if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ such that $\Delta(v_{1},\dots,v_{k})\neq 0$ but $\Delta_{1}(v_{1},\dots,v_{k})=0$.
Note. In Theorem 4.6, we may take $v_{1}=1$.
4.5. Generalized results for $g_{q-1}$
We shall see that most results about $f_{\text{\rm inv}}$ on $\mathbb{F}_{2^{n}}$ can be generalized to $g_{q-1}$ on $\mathbb{F}_{q^{n}}$.
Proposition 4.7.
  • (i)
    The function $g_{q-1}$ is 11st order sum-free.
  • (ii)
    Let $n\geq 2$. Then $g_{q-1}$ is 22nd order sum-free if and only if $n$ is odd.
Proof.
(i) Immediate from Theorem 4.6.
(ii) Let $v_{2}\in\mathbb{F}_{q^{n}}$. Then
$\Delta(1,v_{2})\neq 0\ \Leftrightarrow\ \left|\begin{matrix}1&v_{2}\cr 1&v_{2}^{q}\end{matrix}\right|\neq 0\ \Leftrightarrow\ v_{2}\notin\mathbb{F}_{q},$
and
$\Delta_{1}(1,v_{2})=0\ \Leftrightarrow\ \left|\begin{matrix}1&v_{2}\cr 1&v_{2}^{q^{2}}\end{matrix}\right|=0\ \Leftrightarrow\ v_{2}^{q^{2}}=v_{2}\ \Leftrightarrow\ v_{2}\in\mathbb{F}_{q^{2}}.$
Hence,
there exists $v_{2}\in\mathbb{F}_{q^{2}}$ such that $\Delta(1,v_{2})\neq 0$ but $\Delta_{1}(1,v_{2})=0$
$\Leftrightarrow\mathbb{F}_{q^{2}}\subset\mathbb{F}_{q^{n}}\Leftrightarrow\text{$n$ is even}.$
For $n\geq 2$, define
$\mathcal{K}_{n,q}=\{1\leq k\leq n-1:\text{$g_{q-1}$ is not $k$th order sum-free on $\mathbb{F}_{q^{n}}$}\}.$
When $q=2$, $\mathcal{K}_{n,2}=\mathcal{K}_{n}$, which is defined in equation (1.1).
Proposition 4.8.
If $\gcd(k,n)>1$, then $k\in\mathcal{K}_{n,q}$.
Proof.
Let $d=\gcd(k,n)$ and let $E$ be a $k/d$-dimensional $\mathbb{F}_{q^{d}}$-subspace of $\mathbb{F}_{q^{n}}$. Then $\dim_{\mathbb{F}_{q}}E=k$. Since $d>1$, there exists $a\in\mathbb{F}_{q^{d}}\setminus\mathbb{F}_{q}$. Since $aE=E$, we have
$\sum_{0\neq u\in E}\frac{1}{u^{q-1}}=\sum_{0\neq u\in E}\frac{1}{(au)^{q-1}}=a^{1-q}\sum_{0\neq u\in E}\frac{1}{u^{q-1}},$
where $a^{1-q}\neq 1$. Hence
$\sum_{0\neq u\in E}\frac{1}{u^{q-1}}=0.$
Lemma 4.9.
Let $E$ and $F$ be subspaces of $\mathbb{F}_{q^{n}}$ such that $E\cap F=\{0\}$. Then
(4.11) $\sum_{0\neq u\in E\oplus F}\frac{1}{u^{q-1}}=\sum_{0\neq u\in E}\frac{1}{u^{q-1}}+\Delta(v_{1},\dots,v_{k})^{(q-1)^{2}}\sum_{0\neq v\in L_{E}(F)}\frac{1}{v^{q-1}},$
where $v_{1},\dots,v_{k}$ is a basis of $E$ and $L_{E}(X)=\prod_{u\in E}(X+u)$.
Proof.
We have
(4.12) $\sum_{0\neq u\in E\oplus F}\frac{1}{u^{q-1}}=\sum_{0\neq u\in E}\frac{1}{u^{q-1}}+\sum_{0\neq w\in F}\,\sum_{u\in E}\frac{1}{(w+u)^{q-1}}.$
By equation (4.10),
$\sum_{u\in\langle v_{1},\dots,v_{k}\rangle}\frac{1}{(Y+u)^{q-1}}=\Bigl(\frac{\Delta_{0}(v_{1},\dots,v_{k})}{\Delta(Y,v_{1},\dots,v_{k})}\Bigr)^{q-1}.$
In the above, $\Delta_{0}(v_{1},\dots,v_{k})=\Delta(v_{1},\dots,v_{k})^{q}$, and by equation (4.7),
$\Delta(Y,v_{1},\dots,v_{k})=(-1)^{k}L_{E}(Y)\Delta(v_{1},\dots,v_{k}).$
Hence
(4.13) $\sum_{u\in\langle v_{1},\dots,v_{k}\rangle}\frac{1}{(Y+u)^{q-1}}=\Bigl(\frac{\Delta(v_{1},\dots,v_{k})^{q}}{(-1)^{k}L_{E}(Y)\Delta(v_{1},\dots,v_{k})}\Bigr)^{q-1}=\frac{\Delta(v_{1},\dots,v_{k})^{(q-1)^{2}}}{L_{E}(Y)^{q-1}}.$
Using equation (4.13) in (4.12) gives
$\sum_{0\neq u\in E\oplus F}\frac{1}{u^{q-1}}=\sum_{0\neq u\in E}\frac{1}{u^{q-1}}+\Delta(v_{1},\dots,v_{k})^{(q-1)^{2}}\sum_{0\neq w\in F}\frac{1}{L_{E}(w)^{q-1}}.$
Since $\ker(L_{E}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}})=E$, we have $\ker(L_{E})\cap F=E\cap F=\{0\}$. Hence the restriction $L_{E}|_{F}:F\to L_{E}(F)$ is an isomorphism. Therefore,
$\sum_{0\neq w\in F}\frac{1}{L_{E}(w)^{q-1}}=\sum_{0\neq v\in L_{E}(F)}\frac{1}{v^{q-1}}.$
This completes the proof of the lemma. ∎
Theorem 4.10 ($q$-ary version of [5, Theorem 3]).
Assume $n\geq 2$. If $E$ and $F$ are subspaces of $\mathbb{F}_{q^{n}}$ such that $\mathbb{F}_{q^{n}}=E\oplus F$, then
$\sum_{0\neq u\in E}\frac{1}{u^{q-1}}\neq 0\ \Leftrightarrow\ \sum_{0\neq v\in L_{E}(F)}\frac{1}{v^{q-1}}\neq 0.$
In particular, $k\in\mathcal{K}_{n,q}$ if and only if $n-k\in\mathcal{K}_{n,q}$.
Proof.
In equation (4.11), we have
$\sum_{0\neq u\in E\oplus F}\frac{1}{u^{q-1}}=\sum_{0\neq u\in\mathbb{F}_{2^{n}}}\frac{1}{u^{q-1}}=0$
since $n\geq 2$. The conclusions follow immediately. ∎
For $\mathbb{F}_{q}$-subspaces $E$ and $F$ of $\mathbb{F}_{q^{n}}$, let $EF$ denote the $\mathbb{F}_{q}$-span of $\{xy:x\in E,\ y\in F\}$ and let $F^{\bot}=\{x\in\mathbb{F}_{q^{n}}:\text{\rm Tr}_{q^{n}/q}(xy)=0\ \text{for all}\ y\in F\}$.
Theorem 4.11 ($q$-ary version of §2.2 Result (5)).
If $k,l\in\mathcal{K}_{n,q}$ and $kl<n$, then $k+l\in\mathcal{K}_{n,q}$.
Proof.
Let $E$ and $F$ be $\mathbb{F}_{q}$-subspaces of $\mathbb{F}_{q^{n}}$ such that $\dim_{\mathbb{F}_{q}}E=k$, $\dim_{\mathbb{F}_{q}}F=l$, and $\sum_{0\neq u\in E}1/u^{q-1}=0$, $\sum_{0\neq v\in F}1/v^{q-1}=0$. Since
$\dim_{\mathbb{F}_{q}}\bigl(F\cdot L_{E}(\mathbb{F}_{q^{n}})^{\bot}\bigr)\leq\dim_{\mathbb{F}_{q}}F\cdot\dim_{\mathbb{F}_{q}}E=lk<n,$
there exists $0\neq a\in\mathbb{F}_{q^{n}}$ such that $\text{\rm Tr}_{q^{n}/q}\bigl(aF\cdot L_{E}(\mathbb{F}_{q^{n}})^{\bot}\bigr)=0$, i.e., $aF\subset L_{E}(\mathbb{F}_{q^{n}})$. Since $L_{E}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ induces an isomorphism $\mathbb{F}_{q^{n}}/E\to L_{E}(\mathbb{F}_{q^{n}})$, there exists an $\mathbb{F}_{q}$-subspace $V$ of $\mathbb{F}_{q^{n}}$ such that $V\cap E=\{0\}$, $\dim_{\mathbb{F}_{q}}V=l$ and $L_{E}(V)=aF$. Let $b=\Delta(v_{1},\dots,v_{k})^{(q-1)^{2}}$, where $v_{1},\dots,v_{k}$ is a basis of $E$ over $\mathbb{F}_{q}$. Now by Lemma 4.9,
$\sum_{0\neq u\in E\oplus V}\frac{1}{u^{q-1}}\,$ $=\sum_{0\neq u\in E}\frac{1}{u^{q-1}}+b\sum_{0\neq v\in L_{E}(V)}\frac{1}{v^{q-1}}$
$=b\sum_{0\neq v\in aF}\frac{1}{v^{q-1}}=\frac{b}{a^{q-1}}\sum_{0\neq v\in F}\frac{1}{v^{q-1}}=0.$
Hence $k+l\in\mathcal{K}_{n,q}$. ∎
Theorem 4.12 ($q$-ary version of [7, Theorem 3.1]).
The following two statements are equivalent:
  • (i)
    There exists $x\in\mathbb{F}_{q^{n}}$ such that $\Delta(x,x^{q},x^{q^{2}},\dots,x^{q^{k-1}})\neq 0$ and
    $\Delta_{1}(x,x^{q},x^{q^{2}},\dots,x^{q^{k-1}})=0$.
  • (ii)
    $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$.
Proof.
Let $\sigma$ denote the Frobenius automorphism of $\mathbb{F}_{q^{n}}$ over $\mathbb{F}_{q}$. Note that
$\Delta(x,x^{q},x^{q^{2}},\dots,x^{q^{k-1}})\neq 0$
[…]
and
$\Delta_{1}(x,x^{q},x^{q^{2}},\dots,x^{q^{k-1}})=0$
[…]
(ii) $\Rightarrow$ (i). Let $\alpha\in\mathbb{F}_{q^{n}}$ be a normal element over $\mathbb{F}_{q}$. Write $X^{n}-1=fg$, where $f=X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}$. Let $x=(g(\sigma))(\alpha)$. For each $0\neq h\in\mathbb{F}_{q}[X]$ with $\deg h<k$, $hg\not\equiv 0\pmod{X^{n}-1}$, whence $(h(\sigma))(x)=((hg)(\sigma))(\alpha)\neq 0$. Hence $x,\sigma(x),\dots,\sigma^{k-1}(x)$ are linearly independent over $\mathbb{F}_{q}$.
Since $(f(\sigma))(x)=((fg)(\sigma))(\alpha)=0$, the elements $x,\sigma^{2}(x),\dots,\sigma^{k}(x)$ are linearly dependent over $\mathbb{F}_{q}$.
(i) $\Rightarrow$ (ii). Since $x,\sigma^{2}(x),\dots,\sigma^{k}(x)$ are linearly dependent over $\mathbb{F}_{q}$, there exists $0\neq f=a_{k}X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$ such that $(f(\sigma))(x)=0$. We claim that $a_{k}\neq 0$ and $f\mid X^{n}-1$. Otherwise, $f_{1}:=\text{gcd}(f,X^{n}-1)$ has degree <k<k and $(f_{1}(\sigma))(x)=0$. Then $x,\sigma(x),\dots,\sigma^{k-1}(x)$ are linearly dependent over $\mathbb{F}_{q}$, which is a contradiction. ∎
Corollary 4.13.
If $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{q}[X]$, then $k\in\mathcal{K}_{n,q}$.
4.6. Numerical results
It follows from Propositions 4.7, 4.8 and Theorem 4.10 that
$\mathcal{K}_{n,q}=\begin{cases}\emptyset&\text{if}\ n=2,3,5,\cr\{2\}&\text{if}\ n=4,\cr\{2,3,4\}&\text{if}\ n=6.\end{cases}$
For $q=3,5$ and $7\leq n\leq 11$, the answers to the question whether $g_{q-1}$ is $k$th order sum-free are tabulated in Table 1 (for $q=3$) and Table 2 (for $q=5$). These results are obtained through a combination of the theoretical results in §4.5 and computer searches. It is remarkable that for $q=3,5$ and $n=7$, $g_{q-1}$ is $k$th order sum-free for all $1\leq k\leq 6$; such a phenomenon has not been observed in the binary case.
Table 1. $q=3;\ 7\leq n\leq 11,\ 1\leq k\leq n-1$: Is $g_{q-1}$ $k$th order sum-free?
[…]
7
8
9
10
11
Table 2. $q=5;\ 7\leq n\leq 11,\ 1\leq k\leq n-1$: Is $g_{q-1}$ $k$th order sum-free?
[…]
7
8
9
10
11 ? ?
Appendix. Absolute Irreducibility of $\Theta_{4}$
Recall that
$\Theta_{4}(X_{1},\dots,X_{4})=m_{(2^{3})}+m_{(2^{2},2^{2})}+m_{(2^{2},2,2)}+m_{(2^{2},2,1,1)}+m_{(2,2,2,2)},$
where
$m_{(2^{3})}=X_{1}^{2^{3}}+X_{2}^{2^{3}}+X_{3}^{2^{3}}+X_{4}^{2^{3}},$
$m_{(2^{2},2^{2})}=X_{1}^{2^{2}}X_{2}^{2^{2}}+X_{1}^{2^{2}}X_{3}^{2^{2}}+X_{1}^{2^{2}}X_{4}^{2^{2}}+X_{2}^{2^{2}}X_{3}^{2^{2}}+X_{2}^{2^{2}}X_{4}^{2^{2}}+X_{3}^{2^{2}}X_{4}^{2^{2}},$
$m_{(2^{2},2,2)}=\,$ $X_{2}^{2^{2}}X_{3}^{2}X_{4}^{2}+X_{2}^{2}X_{3}^{2^{2}}X_{4}^{2}+X_{2}^{2}X_{3}^{2}X_{4}^{2^{2}}$
$+X_{1}^{2^{2}}X_{3}^{2}X_{4}^{2}+X_{1}^{2}X_{3}^{2^{2}}X_{4}^{2}+X_{1}^{2}X_{3}^{2}X_{4}^{2^{2}}$
$+X_{1}^{2^{2}}X_{2}^{2}X_{4}^{2}+X_{1}^{2}X_{2}^{2^{2}}X_{4}^{2}+X_{1}^{2}X_{2}^{2}X_{4}^{2^{2}}$
$+X_{1}^{2^{2}}X_{2}^{2}X_{3}^{2}+X_{1}^{2}X_{2}^{2^{2}}X_{3}^{2}+X_{1}^{2}X_{2}^{2}X_{3}^{2^{2}}.$
$m_{(2^{2},2,1,1)}=\,$ $X_{1}^{2^{2}}X_{2}^{2}X_{3}X_{4}+X_{1}^{2^{2}}X_{2}X_{3}^{2}X_{4}+X_{1}^{2^{2}}X_{2}X_{3}X_{4}^{2}$
$+X_{1}^{2}X_{2}^{2^{2}}X_{3}X_{4}+X_{1}X_{2}^{2^{2}}X_{3}^{2}X_{4}+X_{1}X_{2}^{2^{2}}X_{3}X_{4}^{2}$
$+X_{1}^{2}X_{2}X_{3}^{2^{2}}X_{4}+X_{1}X_{2}^{2}X_{3}^{2^{2}}X_{4}+X_{1}X_{2}X_{3}^{2^{2}}X_{4}^{2}$
$+X_{1}^{2}X_{2}X_{3}X_{4}^{2^{2}}+X_{1}X_{2}^{2}X_{3}X_{4}^{2^{2}}+X_{1}X_{2}X_{3}^{2}X_{4}^{2^{2}},$
$m_{(2,2,2,2)}=X_{1}^{2}X_{2}^{2}X_{3}^{2}X_{4}^{2}.$
Theorem A1.
The polynomial $\Theta_{4}(X_{1},\dots,X_{4})$ is irreducible over $\overline{\mathbb{F}}_{2}$.
To prove Theorem A1, we need a couple of lemmas. We have
$\Theta_{4}(X_{1},\dots,X_{4})=s_{4}^{2}+s_{1}(s_{1}s_{2}+s_{3})s_{4}+(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{2}=:G(s_{1},\dots,s_{4}),$
where $s_{i}$ is the $i$th elementary symmetric polynomials in $X_{1},\dots,X_{4}$.
Lemma A2.
The polynomial $G(s_{1},\dots,s_{4})$ is irreducible in $\overline{\mathbb{F}}_{2}[s_{1},\dots,s_{4}]$.
Proof.
We treat $G(s_{1},\dots,s_{4})$ as a polynomial in $s_{4}$. Assume to the contrary that
$G=(s_{4}+A)(s_{4}+B)$
for some $A,B\in\overline{\mathbb{F}}_{2}[s_{1},s_{2},s_{3}]$. Then
$\begin{cases}A+B=s_{1}(s_{1}s_{2}+s_{3}),\cr AB=(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{2}.\end{cases}$
Clearly, $s_{1}^{4}+s_{2}^{2}+s_{1}s_{3}$ is irreducible in $\overline{\mathbb{F}}_{2}[s_{1},s_{2},s_{3}]$. Hence $A=a(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{i}$ and $B=b(s_{1}^{4}+s_{2}^{2}+s_{1}s_{3})^{2-i}$ for some $a,b\in\overline{\mathbb{F}}_{2}^{*}$ and $i\in\{0,1,2\}$. In any case, $A+B$ does not contain any monomial of degree $3$. Thus $A+B\neq s_{1}(s_{1}s_{2}+s_{3})$, which is a contradiction. ∎
Lemma A2 implies that $\Theta_{4}(X_{1},\dots,X_{4})$ does not have a proper factor that is symmetric in $X_{1},\dots,X_{4}$.
We have
$\Theta_{4}(X_{1},X_{2},X_{3},0)=H(X_{1},X_{2},X_{3})^{2},$
where
$H(X_{1},X_{2},X_{3})=X_{1}^{4}+X_{2}^{4}+X_{3}^{4}+X_{1}^{2}X_{2}^{2}+X_{1}^{2}X_{3}^{2}+X_{2}^{2}X_{3}^{2}+X_{1}X_{2}X_{3}(X_{1}+X_{2}+X_{3}).$
Lemma A3.
The polynomial $H(X_{1},X_{2},X_{3})$ is irreducible over $\overline{\mathbb{F}}_{2}$.
Proof.
We treat $H$ as a polynomial in $X_{3}$:
$H=X_{3}^{4}+(X_{1}^{2}+X_{2}^{2}+X_{1}X_{2})X_{3}^{2}+X_{1}X_{2}(X_{1}+X_{2})X_{3}+(X_{1}^{2}+X_{2}^{2}+X_{1}X_{2})^{2},$
where
$X_{1}^{2}+X_{2}^{2}+X_{1}X_{2}=(X_{1}+\alpha X_{2})(X_{1}+\alpha^{-1}X_{2}),\quad\alpha\in\mathbb{F}_{2^{2}}\setminus\mathbb{F}_{2}.$
First assume that $H$ has a factor $X_{3}+A$, where $A\in\overline{\mathbb{F}}_{2}[X_{1},X_{2}]$ is homogeneous of degree $1$. Then $A\mid(X_{1}+\alpha X_{2})(X_{1}+\alpha^{-1}X_{2})$, say $A=c(X_{1}+\alpha X_{2})$, where $c\in\overline{\mathbb{F}}_{2}^{*}$. Then we have
$0=\,$ $H(X_{1},X_{2},c(X_{1}+\alpha X_{2}))$
$=\,$ $(1+c+c^{2})^{2}X_{1}^{4}+c(1+c)X_{1}^{3}X_{2}+(1+c)(1+\alpha c)X_{1}^{2}X_{2}^{2}$
$+c(\alpha+c+\alpha c)X_{1}X_{2}^{3}+(1+c)^{2}(1+\alpha c^{2})X_{2}^{4}.$
It follows that
$\begin{cases}1+c+c^{2}=0,\cr c(1+c)=0,\end{cases}$
which is impossible.
Now assume that $H$ has an absolute irreducible factor $h=X_{3}^{2}+AX_{3}+B$, where $A,B\in\overline{\mathbb{F}}_{2}[X_{1},X_{2}]$. It follows that $B=c(X_{1}+\alpha X_{2})^{2}$ or $c(X_{1}^{2}+X_{2}^{2}+X_{1}X_{2})$, where $c\in\overline{\mathbb{F}}_{2}^{*}$.
Case 1. Assume that $B=c(X_{1}+\alpha X_{2})^{2}=c(X_{1}^{2}+\alpha^{-1}X_{2}^{2})$. Then $X_{3}^{2}+A(X_{2},X_{1})X_{3}+B(X_{2},X_{1})$ is a different irreducible factor of $H$ (different since $B(X_{2},X_{1})\neq B(X_{1},X_{2})$). Thus
$H=(X_{3}^{2}+A(X_{1},X_{2})X_{3}+B(X_{1},X_{2}))(X_{3}^{2}+A(X_{2},X_{1})X_{3}+B(X_{2},X_{1})).$
Comparing the coefficients gives
$A(X_{1},X_{2})+A(X_{2},X_{1})=0,$
$B(X_{1},X_{2})+B(X_{2},X_{1})+A(X_{1},X_{2})A(X_{2},X_{1})=X_{1}^{2}+X_{2}^{2}+X_{1}X_{2}.$
It follows that
$c\alpha(X_{1}^{2}+X_{2}^{2})+A(X_{1},X_{2})^{2}=X_{1}^{2}+X_{2}^{2}+X_{1}X_{2}.$
This is impossible since $X_{1}X_{2}$ does not appear at the left hand side.
Case 2. Assume that $B=c(X_{1}^{2}+X_{2}^{2}+X_{1}X_{2})$.
Case 2.1. Assume that $A(X_{1},X_{2})=A(X_{2},X_{1})$. Then $A(X_{1},X_{2})=b(X_{1}+X_{2})$ for some $b\in\overline{\mathbb{F}}_{2}$, and hence
$h=X_{3}^{2}+b(X_{1}+X_{2})X_{3}+c(X_{1}^{2}+X_{2}^{2}+X_{1}X_{2}).$
However,
$H\equiv\,$ $X_{3}\left(b^{3}X_{1}^{3}+b^{3}X_{1}^{2}X_{2}+b^{3}X_{1}X_{2}^{2}+b^{3}X_{2}^{3}+bX_{1}^{3}+bX_{2}^{3}+X_{1}^{2}X_{2}+X_{1}X_{2}^{2}\right)$
$+b^{2}cX_{1}^{4}+b^{2}cX_{1}^{3}X_{2}+b^{2}cX_{1}X_{2}^{3}+b^{2}cX_{2}^{4}+c^{2}X_{1}^{4}+c^{2}X_{1}^{2}X_{2}^{2}+c^{2}X_{2}^{4}$
$+cX_{1}^{4}+cX_{1}^{2}X_{2}^{2}+cX_{2}^{4}+X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}$
$\not\equiv\,$ $0\pmod{h}$,
which is a contradiction.
Case 2.2. Assume that $A(X_{1},X_{2})\neq A(X_{2},X_{1})$. Then
$h(X_{2},X_{1},X_{3})=X_{3}^{2}+A(X_{2},X_{1})X_{3}+B(X_{2},X_{1})$
is a different irreducible factor of $H$. Thus
$H=(X_{3}^{2}+A(X_{1},X_{2})X_{3}+B(X_{1},X_{2}))(X_{3}^{2}+A(X_{2},X_{1})X_{3}+B(X_{2},X_{1})).$
Comparing the coefficients of $X_{3}^{3}$ gives $0=A(X_{1},X_{2})+A(X_{2},X_{1})$, which is a contradiction. ∎
Proof of Theorem A1.
Assume to the contrary that $\Theta_{4}$ has an absolutely irreducible factor $f$ of degree $d$ with $0<d<8$. Treating $f$ as a polynomial in $X_{4}$ over $\overline{\mathbb{F}}_{2}[X_{1},X_{2},X_{3}]$, we may write
$f=X_{4}^{d}+A_{d-1}X_{4}^{d-1}+\cdots+A_{0},$
where $A_{i}\in\overline{\mathbb{F}}_{2}[X_{1},X_{2},X_{3}]$ is homogeneous of degree did-i. Then $A_{0}\mid\Theta_{4}(X_{1},X_{2},X_{3},0)=H^{2}$. Since $H$ is absolutely irreducible (Lemma A3) and $\deg H=4$, we must have $d=4$ and $A_{0}=cH$ for some $c\in\overline{\mathbb{F}}_{2}$, so
$f=X_{4}^{4}+A_{3}X_{4}^{3}+A_{2}X_{4}^{2}+A_{1}X_{4}+cH.$
The symmetric group $\mathfrak{S}_{3}$ acts on $\overline{\mathbb{F}}_{2}[X_{1},\cdots,X_{4}]$ by permuting $X_{1},X_{2},X_{3}$. We claim that $f$ is symmetric in $X_{1},X_{2},X_{3}$. Otherwise, there exists $\sigma\in\mathfrak{S}_{3}$ such that
$\sigma(f)=X_{4}^{4}+\sigma(A_{3})X_{4}^{3}+\sigma(A_{2})X_{4}^{2}+\sigma(A_{1})X_{4}+cH$
is a different irreducible factor of $\Theta_{4}$. It follows that
$\Theta_{4}=f\sigma(f)=X_{4}^{8}+\cdots+(A_{1}+\sigma(A_{1}))cHX_{4}+c^{2}H^{2}.$
However, the coefficient of $X_{4}$ in $\Theta_{4}$ is
$X_{1}X_{2}X_{3}(X_{1}+X_{2})(X_{1}+X_{3})(X_{2}+X_{3})(X_{1}+X_{2}+X_{3}),$
which is not divisible by $H$. This is a contradiction, hence the claim is proved.
We may also treat $f$ as a polynomial in $X_{1}$ over $\overline{\mathbb{F}}_{2}[X_{2},X_{3},X_{4}]$. By the same argument, $f$ is also symmetric in $X_{2},X_{3},X_{4}$. It follows that $f$ is symmetric in $X_{1},\dots,X_{4}$. However, this is impossible by Lemma A2.
The proof of Theorem A1 is now complete. ∎
We suspect that in general, $\Theta_{k}$ is absolutely irreducible for $k\geq 3$, although the proof for the case $k=4$ given here is already quite technical.
Acknowledgments
This work was supported by NSF REU Grant 2244488.